# LLM4SBR: A Lightweight and Effective Framework for Integrating Large Language Models in Session-based Recommendation

### v1

@@ 4.1.2. Evaluation metrics
In terms of the evaluation indicators used in the experiment, We chose the most commonly used ones in SBR tasks: Precision (P) @$K$ and Mean reciprocal rank (MRR) @$K$. After referring to the classic work (Wang et al., 2020; Xia et al., 2021) in recent years, we set the length of the candidate set @$K$ to $5$, $10$, and $20$, which is the most meaningful for comparison.
4.1.3. Parameter Settings
All experiments were conducted on NVIDIA A100 GPUs. For fairness in performance comparison, the optimizer used throughout the experiments was unified as Adam with a learning rate of $0.001$, decayed by $0.1$ every three epochs, and an $L2$ penalty set to $10^{-5}$. For the SBR model involved in the experiments, the batch size is $100$ and the dimension size is $100$. $\tau$ is set to 0.1. We initially set the hyperparameter $f$ in the intent localization module to $5$, and subsequent hyperparameter experiments 4.4 will discuss the optimal value. We followed the optimal parameter settings as published in their paper for the remaining parameters.
4.2. Performance Experiment and Analysis
@@
4.2.1. Backbone
To validate the effectiveness of the framework, we selected four classic models from SBR to replace the SBR model in the framework and compared the performance between each pair. The introduction of the SBR models is as follows:
@@ 4.2.1. Backbone
SR-GNN 6.30 10.02 14.86 3.18 3.70 3.99 4.29 8.64 13.01 2.16 3.09 3.19
LLM4SBR(SR-GNN) 7.58 11.29 16.30 4.34 4.62 5.00 7.38 11.52 17.54 4.06 4.55 5.26
SR-GNN Improv. 20.31% 12.67% 9.69% 36.47% 24.86% 25.31% 72.02% 33.33% 34.82% 87.96% 47.25% 64.89%
@@
The comparison results of the overall performance experiments are shown in Table 2. We record the performance with K set to 5, 10, 20. It is worth noting that in the evaluation system of RS, smaller $K$ values are more significant. From the results displayed in Table 2, we draw the following observations:
@@
  • LLM4SBR significantly improves backbone performance. In the models enhanced through the LLM framework, all demonstrate performance improvement. This confirms that the text representations derived from LLM inference contain rich and valuable information, which can greatly help the SBR model understand the potential intention of the conversation data.
@@
  • LLM4SBR has a greater improvement for smaller $K$ values. For example, LLM4SBR (TAGNN) improved the P@5 index of the two data sets by 27.28% and 107.5% respectively. We believe this is due to the semantic enhancement achieved by LLM4SBR during the intent localization stage, where it utilizes $f$ similar semantic items. Consequently, it results in more accurate predictions for the top few items in the predicted candidate set. We also observe slight decreases in performance for $S^{2}$-DHCN and GCE-GNN on a few metrics ($P@20$ and $MRR@20$) after integrating with the framework. We posit that when the original SBR model already effectively models the data, enhancing the inference information through the intent localization module may introduce noise. Compared to the improvement magnitude, the decrease is very slight. Moreover, since noise issues can be effectively controlled by adjusting the hyperparameter $f$ in the intent localization module, the negative impact can be almost negligible.
@@
  • LLM4SBR can compensate for poor modeling caused by a lack of interactive information. GCE-GNN captures effective information at both the item and session levels by constructing global graphs and session graphs simultaneously, due to the model’s complex computations, in scenarios with limited data volume, it becomes challenging for this model to learn effective session representations. LLM4SBR (GCE-GNN) showed the greatest improvement, especially on the ML-1M dataset, P@5, P@10, and P@20 increased by 37.59%, 96.2%, and 128.54% respectively. We attribute this to the effective text information obtained from LLM inference, which compensates for the information scarcity in GCE-GNN’s session modeling, allowing it to achieve better performance. In addition, The architectures of SR-GNN and TAGNN are based on directed session graphs, utilizing GNN to capture complex transition relationships between items. However, limited by the number of layers in GNNs, both of these models struggle to effectively capture useful information from long-term items. After adding LLM4SBR, both of the above two models have achieved substantial performance improvements. Inference information from a long-term perspective solves the problem of insufficient capture of long-term dependencies in the model.
@@
4.3. Ablation Experiment and Analysis
(a) $P@K$ on Beauty
(b) $MRR@K$ on Beauty
(c) $P@K$ on ML-1M
(d) $MRR@K$ on ML-1M
Figure 4. Inference perspective ablation experiment results
@@ 4.3. Ablation Experiment and Analysis
To examine the necessity and relative importance of the long-term and short-term inference perspectives, we designed two variants: LLM4SBR w/o Long and LLM4SBR w/o Short. LLM4SBR w/o Long indicates inference without considering the long-term perspective, retaining only the short-term perspective. Conversely, LLM4SBR w/o Short retains only the long-term perspective and removes the short-term perspective during inference. We compared the performance of these two variants with the whole performance and visualized the comparison as a bar chart to clearly illustrate the differences between them.
@@
  • Both long-term and short-term perspectives are necessary. Because the whole framework represented by the blue column in the figure shows the best performance on both datasets. This demonstrates the need to leverage LLM for separate perspective inference, removing any perspective will result in a performance loss.
@@
  • The contribution of long-term and short-term perspective inference varies across the two datasets. In Beauty, the framework relies more on the information provided by the short-term perspective, as discarding the inference results of the short-term perspective would lead to a greater performance drop. Conversely, in Ml-1M, it’s the opposite; the framework relies more on the inference results of the long-term perspective. Through discussion and analysis, we attribute this performance difference to the length of the dataset sessions. Session intent in short sequences is usually relatively stable, and the intent is mainly reflected in the last few clicks. This underscores the increased importance of accurately modeling short-term interests in short-session scenarios. However, as the session length increases, the session intent is influenced by various factors, thereby increasing the importance of long-term dependency relationships within the session. Finally, we believe that considering the inference results of multiple perspectives simultaneously can enhance the stability of the framework’s performance, making it adaptable to datasets with varying session lengths.
4.4. Hyperparameter Experiment and Analysis
[…]
Figure 5. Hyperparameter experimental results of different f settings of the Intent positioning module on ML-1M
@@ 4.4. Hyperparameter Experiment and Analysis
In this section, we discuss the hyperparameter $f$ set within the intent localization module. This hyperparameter is designed to eliminate hallucination and enhance semantics in the preliminary inference results of LLM, using a candidate set of items with similar semantics. The hyperparameter $f$ is configured to control the range of selecting items with similar semantics. The value of $f$ is set to 0, 1, 3, and 5, and we discuss four scenarios accordingly: (1) directly utilizing the inference results of LLM, (2) eliminating hallucination using the most similar item, (3) eliminating hallucination and enhancing semantics using the Top-$3$ most similar items, and (4) eliminating hallucination and enhancing semantics using the Top-$5$ most similar items.
@@
The experimental results are shown in Figure 5. Firstly, across all three plots, although the optimal hyperparameter values differ for each plot, we can see that the performance is consistently the worst when $f=0$. We believe this is logical and demonstrates the necessity of the intent localization module in the framework. If the results of LLM inference are not processed, hallucinations occurring in some session data may lead to a decrease in the overall framework performance. Taking a closer look at the local details, in Figures 5 and 5, the performance peaks when $f=1$, with $P@5$ and $P@10$ being 7.69 and 11.48 respectively. However, in Figure 5, a notable peak is observed when $f=3$, achieving the best performance with $P@20$ is 18.56. When the value of $K$ is relatively small $(K=5,K=10)$, the performance is best when $f=1$. We believe this is because utilizing multiple similar items for semantic enhancement of intent may introduce noise, thereby leading to a slight performance decrease. As the value of $K$ increases, the performance of the $f=3$ and $f=5$ becomes similar, both surpass $f=1$. This suggests that introducing multiple similarly named items appropriately can increase the diversity of candidate items while enhancing performance.
In summary, values of $f$ ranging from $1$ to $5$ are all effective. Depending on the requirements for different values of $K$, selecting different values of $f$ can better leverage the module’s effectiveness.
4.5. Space Occupation Experiment and Analysis
Table 3. The result of space occupation experiment
Model Occupies video memory (M)

### v2

@@ 5.1.2. Evaluation metrics
In terms of the evaluation indicators used in the experiment, We chose the most commonly used ones in SBR tasks: Precision (P) @$K$, Mean reciprocal rank (MRR) @$K$ and Normalized Discounted cumulative gain (NDCG) @$K$. After referring to the classic work (Wu et al., 2019; Wang et al., 2020; Xia et al., 2021; Huang et al., 2021)  [^4] The calculation method of $P@20$ and $MRR@20$ refer to the (Wu et al., 2019), and $NDCG@K$ refers to (Huang et al., 2021). in recent years, we set the length of the candidate set @$K$ to $5$, $10$, and $20$, which is the most meaningful for comparison.
5.1.3. Implementation details and Hyper-parameter settings
All experiments were conducted on NVIDIA A100 GPUs. In the aspect of selecting large models, we have chosen the Qwen-7B-Chat  [^5] https://github.com/QwenLM/Qwen model as the inference model after careful consideration of LLM’s inference capability, adaptability to both Chinese and English languages and model parameter count. For fairness in performance comparison, the optimizer used throughout the experiments was unified as Adam with a learning rate of $0.001$, decayed by $0.1$ every three epochs, and an $L2$ penalty set to $10^{-5}$. For the SBR model involved in the experiments, the batch size is $100$ and the dimension size is $100$. $\tau$ is set to 0.1. We initially set the hyperparameter $r$ in the intent localization module to $5$, and subsequent hyperparameter experiments 5.4 will discuss the optimal value. We followed the optimal parameter settings as published in their paper for the remaining parameters.
5.2. Performance Experiment and Analysis (RQ1)
@@
5.2.1. Backbones
To validate the effectiveness of the framework, we carefully selected six classic models in the field of Session-based Recommendation/Sequential Recommendation, with the SBR model as its core component. Note that here we have not chosen to compare performance with a model that incorporates LLM. This is because there is very little work on SBR incorporating LLM, and no open-source code exists. Second, the architecture of our work is different from other work, focusing on enhancements to the traditional SBR model rather than from the view of replacing SBR with LLM. Among them, the GRU4Rec and SASRec models are particularly good at mining sequences’ temporal dynamics and patterns. In contrast, SRGNN and TAGNN focus more on revealing the transfer or transformation relationships between items. The GCE-GNN and DHCN models, based on effectively capturing item transformation relationships, further incorporate the concept of multi-graph structure and introduce the use of inter-session information to enhance the model’s ability to understand complex interaction patterns. In the experimental section, we assess the performance of each model both before and after enhancement with the framework and systematically compare and analyze the performance disparities among the enhanced models.
The introduction of the SBR models is as follows:
  • GRU4Rec (Hidasi et al., 2015) utilizes GRU to learn dependencies between sequences to predict the next likely item of interest.
  • SASRec (Kang and McAuley, 2018) uses a self-attention mechanism to solve the long-term dependency problem existing in traditional sequence models.
@@ 5.2.1. Backbones
GRU4Rec 4.99 8.18 12.71 2.71 3.06 3.23 5.48 9.52 14.72 2.98 3.40 3.57
LLM4SBR(GRU4Rec) 7.50 11.68 16.74 4.51 4.89 4.99 6.66 10.91 17.20 3.97 4.38 4.52
GRU4Rec Improv. 50.30% 42.96% 31.68% 66.42% 60.32% 54.60% 21.53% 14.70% 16.88% 33.22% 28.71% 26.72%
SASRec 3.93 6.40 10.11 2.09 2.37 2.54 2.07 3.67 6.47 1.07 1.27 1.42
LLM4SBR(SASRec) 5.29 8.50 13.20 2.98 3.33 3.49 3.44 6.06 10.05 1.72 2.02 2.21
SASRec Improv. 34.6% 32.81% 30.56% 42.58% 40.51% 37.94% 66.18% 65.12% 55.33% 60.74% 59.06% 55.63%
SR-GNN 6.23 10.06 15.06 3.45 3.84 3.99 4.42 7.53 12.27 2.44 2.80 2.98
LLM4SBR(SR-GNN) 7.78 11.73 16.98 4.71 5.03 5.12 8.13 11.80 18.35 4.85 5.16 5.27
SR-GNN Improv. 24.87% 16.60% 12.74% 36.52% 30.98% 28.32% 83.93% 56.70% 49.55% 98.77% 84.28% 76.84%
@@
The comparison results of the overall performance experiments are shown in Table 3. We record the performance with K set to 5, 10, 20. It is worth noting that smaller $K$ values are more significant in the evaluation system of RS. From the results displayed in Table 3, we draw the following observations:
@@
  • LLM4SBR significantly improves backbone performance. In the models enhanced through the LLM framework, both sequence and graph-structured models show significant performance improvements. For example, the $P@5$ of SASRec and SR-GNN on the Ml-1M dataset have increased by 66.18% and 83.93%, respectively. This confirms that the text representations derived from LLM inference contain rich and valuable information, which can greatly help the SBR model understand the potential intention of the conversation data.
@@
  • LLM4SBR has a greater improvement for smaller $K$ values. Almost all backbone models achieve larger improvements at smaller $K$ values when combined with the framework. For example, LLM4SBR (TAGNN) improved the $P@5$ index of the two data sets by 27.28% and 107.5% respectively. We believe this is due to the semantic enhancement achieved by LLM4SBR during the intent localization stage, where it utilizes $r$ similar semantic items. Consequently, it results in more accurate predictions for the top few items in the predicted candidate set. We also observe slight decreases in performance for $S^{2}$-DHCN and GCE-GNN on a few metrics ($P@20$ and $MRR@20$) after integrating with the framework. We posit that when the original SBR model already effectively models the data, enhancing the inference information through the intent localization module may introduce noise. Compared to the improvement magnitude, the decrease is very slight. Moreover, since noise issues can be effectively controlled by adjusting the hyperparameter $r$ in the intent localization module, the negative impact can be almost negligible.
@@
  • The LLM4SBR framework can effectively alleviate modeling defects caused by scene mismatch and data sparseness and fully activate the inherent potential of the model. Due to the short sequence characteristics of session data, although SASRec is good at capturing long-term dependencies, this becomes a limitation in short sequence scenarios, resulting in poor performance. However, after being integrated into the LLM4SBR framework, its shortcomings in short-sequence processing have been significantly compensated, and its performance has been greatly improved. This achievement not only reflects the efficiency of framework optimization and collaborative model capabilities but also proves that framework integration can effectively enhance the model potential in specific scenarios. GCE-GNN captures effective information at both the item and session levels by constructing global graphs and session graphs simultaneously, due to the model’s complex computations, in scenarios with limited data volume, it becomes challenging for this model to learn effective session representations. LLM4SBR (GCE-GNN) showed the greatest improvement, especially on the Ml-1M dataset, $P@5$, $P@10$, and $P@20$ increased by 37.59%, 96.2%, and 128.54% respectively. We attribute this to the effective text information obtained from LLM inference, which compensates for the information scarcity in GCE-GNN’s session modeling, allowing it to achieve better performance.
@@
5.2.2. Comparison with SBR Models Combining LLMs
Figure 5. Performance Comparison with SBR Models Combining LLMs.
In this subsection, we further compare the performance of the LLM4SBR framework with other frameworks that combined with LLM to explore the enhancement capability of LLM in LLM4SBR for SBR models. Note that here we use the same backbone, SRGNN. the specific model is described below:
  • LLM-Infer directly predicts the next click using LLM inference of long-term and short-term interests after intent localization.
  • LLMSeqSim(Harte et al., 2023) retrieves semantically rich embeddings for each item in the session from existing LLMs and computes aggregated session embeddings to recommend products with similar embeddings.
  • LLM2SRGNN (Boz et al., 2024) [^6] We used the LLM2Sequential framework but replaced the SR model backbone with SR-GNN to ensure the fairness of the experiments. initializes the SBR model with the item embeddings obtained from LLM.
In Figure 5, we show the original SR-GNN model in yellow. It can be clearly seen from the figure that LLM-Infer represented in gray and LLMSeqSim represented in blue are significantly lower in performance than SR-GNN, which effectively models collaborative information. This observation demonstrates that although LLMs perform well on many NLP tasks, they are not suitable for direct application in recommendation tasks. At the same time, the other two frameworks (LLM2SRGNN and LLM4SBR) that use LLM to enhance the collaborative model in the figure have achieved significant performance improvements compared to the backbone (SR-GNN). This experimental result confirms that combining LLM with the collaborative model is a wise direction. LLM can provide rich semantic modality information for the collaborative model that is limited by data sparsity. It is particularly noteworthy that LLM4SBR achieved the best results among all models, and achieved a significant performance improvement compared to the second-best SR-GNN, which proves the rationality and effectiveness of our framework design. We conclude that it is not enough to simply introduce knowledge from LLM, but also to effectively integrate information from different modalities and improve the effective utilization of information to achieve more accurate recommendation results.
5.3. Ablation Study (RQ2)
5.3.1. The impact of LLMSBR framework integration on performance
Table 4. The ablation results of LLM and SBR models in the LLM4SBR framework.
Dataset Beauty Ml-1M
Model w/o SBR w/o LLM w/o AU LLM4SBR w/o SBR w/o LLM w/o AU LLM4SBR
$P@5$ 1.92 6.23 7.12 7.78 1.64 4.42 6.31 8.13
$P@10$ 3.44 10.06 10.75 11.73 2.90 7.53 10.07 11.80
$P@20$ 5.76 15.06 15.74 16.98 5.38 12.27 15.11 18.35
$MRR@5$ 0.95 3.45 4.39 4.71 0.83 2.44 3.85 4.85
$MRR@10$ 1.14 3.84 4.69 5.03 0.99 2.80 4.20 5.16
$MRR@20$ 1.27 3.99 4.79 5.12 1.13 2.98 4.34 5.27
$NDCG@5$ 1.21 4.27 5.21 5.64 1.04 3.02 4.54 5.89
$NDCG@10$ 1.69 5.38 6.22 6.67 1.44 3.98 5.60 6.90
$NDCG@20$ 2.23 6.39 7.16 7.64 2.03 4.97 6.65 8.10
To test the integration effect of the LLM4SBR framework, we design three variants as follows:
  • LLM4SBR w/o SBR - using the long-term and short-term interest results of LLM inference after intention localization for direct prediction.
  • LLM4SBR w/o LLM - using the SBR model (here we use SR-GNN) for direct prediction.
  • LLM4SBR w/o AU - removing the representation alignment operation, and directly fuses the LLM inference representation with the long and short-term session representation constructed by SBR.
We added $NDCG@K$ as an evaluation criterion on top of the $P@K$ and $MRR@K$ evaluation metrics. Then we compared the performance of these three variants with the full version of LLM4SBR on the Beauty and Ml-1M datasets. The experimental results are shown in Table 4, where we marked the best-performing model in bold.
By observing and analyzing the results in Table 4, we draw the following conclusions:
  • LLM inference results are ineffective when directly used for recommendation tasks. In the experimental results shown in Table 4, the LLM4SBR w/o SBR variant performed poorly, ranking last across all evaluation metrics. This finding is consistent with previous research (Bao et al., 2023b; Gao et al., 2023), which suggests that there is a significant difference between pre-trained LLMs and the tasks required for the RS, leading to their limited effectiveness when directly applied to SBR scenarios. RSs require models to understand user historical behavior patterns, capture real-time changes in user interests, and effectively match associations within a high-dimensional item space. In contrast, LLMs primarily focus on coherence and meaning expression within language structures, lacking the in-depth understanding of domain-specific knowledge and context required for the recommendation field. Therefore, the core of this study is the innovative integration strategy, which aims to combine the powerful language understanding of LLM with the targetedness of traditional recommendation algorithms to build a more efficient and intelligent RS.
  • The LLM4SBR framework effectively combines the advantages of SBR and LLM. The SBR model using the LLM4SBR framework has the best performance in Table 4, in which the $NDCG@K$ metrics are significantly enhanced. This improvement demonstrates the advantages of LLM4SBR in accurately capturing user interests and proves that the LLM4SBR framework can skillfully integrate the expertise of LLM and SBR to achieve high efficiency in collaborative modeling. The second best performance of all metrics is the LLM4SBR w/o AU, which firstly proves the validity of the operation of split-view representation alignment in the framework, which is more scientifically sound to align and uniform the representations of different modalities in separate views before fusing them, than to directly fuse the different modal representations, and secondly, this variant proves once again that the results of the LLM inference can indeed significantly enhance the performance of the SBR, even without the representation alignment operation, it also performs better than the modeling approach using unimodal information. Through the LLM4SBR framework, the SBR model can not only deeply understand the text context and extract more detailed and comprehensive user preferences, but also effectively handle sequence dependencies in conversational recommendation scenarios, thereby generating more personalized recommendation results.
5.3.2. The influence of LLM inference results from different views
(a) $P@K$ on Beauty.
(b) $MRR@K$ on Beauty.
(c) $P@K$ on Ml-1M.
(d) $MRR@K$ on Ml-1M.
Figure 6. The ablation results from different inference views.
(a) The result of $P@5$ and $MRR@5$
(b) The result of $P@10$ and $MRR@10$
(c) The result of $P@20$ and $MRR@20$
Figure 7. Hyperparameter experimental results of different $r$ settings of the intent localization module on Ml-1M.
To examine the necessity and relative importance of the long-term and short-term inference views, we designed three variants as follows:
  • LLM4SBR w/o Long indicates inference without considering the long-term view, retaining only the short-term view.
  • LLM4SBR w/o Short retains only the long-term view and removes the short-term view during inference.
  • LLM4SBR-Normal removes the view-limiting qualifier in prompt, and LLM directly infers the overall interest, which is then aligned and uniformed with the comprehensive session embedding learned by the traditional SBR model.
@@ 5.3.2. The influence of LLM inference results from different views
We compared the performance of these three variants with the whole performance and visualized the comparison as a bar chart to clearly illustrate the differences between them.
@@
  • Combining prompts from different views can improve the accuracy of LLM in inferring user interests, both long-term and short-term views are necessary. Because the whole framework represented by the blue column in the figure shows the best performance on both datasets. For example, on the Beauty, the $P@5$, $P@10$, and $P@20$ of LLM4SBR are 7.78, 11.73, and 16.98, respectively, while the corresponding w/o Long and w/o Short variants show varying degrees of decreases in all three of the above metrics($P@5$, $P@10$, $P@20$), with values of 7.32, 10.80, and 15.98, as well as 6.92, 10.59, and 15.90. This justifies the necessity for multi-view inference, where deleting any view results in a performance loss. In addition, the second best performing model is LLM-Normal, with $P@5$, $P@10$, and $P@20$ metrics of 7.50, 11.38, and 16.72 on the Beauty dataset, which indicates that inference from any single view is inferior to direct comprehensive inference, but combining prompts from different views can improve the accuracy of LLM’s inference of user interests and effectively improve the recommendation metrics.
@@
  • The contribution of long-term and short-term view inference varies across the two datasets. Specifically, on the Beauty dataset, the $P@5$ of LLM w/o Long and LLM w/o Short are 7.32 and 6.92 respectively, and the former performs better than the latter; on the Ml-1M dataset, the $P@5$ of LLM w/o Long and LLM w/o Short are 6.90 and 7.22 respectively, and the former performs worse than the latter. In Beauty, the framework relies more on the information provided by the short-term view, as discarding the inference results of the short-term view would lead to a greater performance drop. Conversely, in Ml-1M, it’s the opposite; the framework relies more on the inference results of the long-term view. Through discussion and analysis, we attribute this performance difference to the length of the dataset sessions. Session intent in short sequences is usually relatively stable, and the intent is mainly reflected in the last few clicks. This underscores the increased importance of accurately modeling short-term interests in short-session scenarios. However, as the session length increases, the session intent is influenced by various factors, thereby increasing the importance of long-term dependency relationships within the session. Finally, we believe that simultaneously considering the inference results of multiple views can enhance the stability of the framework’s performance, making it adaptable to datasets with varying session lengths.
  • The contribution of view splitting is to enhance the directionality of LLM. The performance of LLM4SBR-Normal is the second best on both datasets, and the performance degradation is not significant. We analyzed the experimental results and believe that there are two main reasons. First, LLM itself has strong reasoning ability. Without clear guidance, LLM can combine global context information and summarize the user’s overall interests on its own. Second, the long-term and short-term view reasoning is mainly used to enhance LLM’s fine-grained understanding of user interests, rather than completely changing its reasoning method. By introducing different view qualifiers in the prompts, the framework can more accurately guide LLM to infer users’ short-term preferences and long-term interests, supplement the missing long-tail items, and thus give full play to its reasoning ability.
In conclusion, within the LLM4SBR framework, each module is indispensable. The LLM4SBR framework successfully integrates the strengths of LLM and SBR models by leveraging LLM for multi-view interest inference and aligning semantic and collaborative information from corresponding views, significantly enhancing recommendation performance.
5.4. Hyperparameter Experiment and Analysis (RQ3)
In this section, we discuss the hyperparameter $r$ set within the intent localization module. This hyperparameter is designed to alleviate hallucination and enhance semantics in the preliminary inference results of LLM, using a candidate set of items with similar semantics. The reason we chose to conduct this experiment on the Ml-1M dataset is that this dataset does not use the data enhancement method of sequence segmentation, which is more reflective of the actual effect of precise localization and semantic enhancement in real scenarios using items with similar semantics.
@@ 5.4. Hyperparameter Experiment and Analysis (RQ3)
The hyperparameter $r$ is configured to control the range of selecting items with similar semantics. The value of $r$ is set to 0, 1, 3, and 5, and we discuss four scenarios accordingly: (1) directly utilizing the inference results of LLM; (2) using the most similar 1 items to alleviate hallucination and enhance semantics; (3) using the most similar 3 items; (4) using the most similar 5 items.
@@
The experimental results are shown in Figure 7. Firstly, across all three subfigures, although the optimal hyperparameter values differ for each subfigure, it can be seen that in most cases, the performance is the worst when $r=0$. We believe this is logical and demonstrates the necessity of the intent localization module in the framework. If the results of LLM inference are not processed, hallucinations occurring in some session data may decrease the overall framework performance. Moving on to each of the figures, in Figure 7(a), the performance enhances as $r$ increases, peaking at $P@5$ and $MRR@5$ when $r=5$, with $P@5$ and $MRR@5$ at 8.13 and 4.85, respectively. In Figures 7(b) and 7(c), the peaks are achieved when $r=1$ or $r=3$, respectively. However, $r$ has little effect on those metrics. We believe that this is because the $r$-value range set by the framework is small (no more than 5), so the intent Localization module only screens the top few items in text similarity for semantic enhancement. Therefore, when evaluating the Top-$K$ metrics, it has a positive impact on smaller $K$ values but has no obvious impact on larger $K$ values.
In summary, $r$ values from $1$ to $5$ are valid. A smaller value of $r$ can play the efficacy of pinpointing, and a larger value of $r$ can improve the accuracy of the whole recommendation list through semantic enhancement. Depending on the actual situation, choosing different $r$ values can better utilize the effectiveness of the module.
5.5. Model training space occupancy experiment and Time complexity analysis (RQ4)
Table 5. The results of GPU usage and training time of model training on the Beauty and Ml-1M datasets.